The weight of an object is the same on two different planets. The mass of planet A is only 60 percent that of planet B. Find the ratio rA/rB of the radii of the planets.
Answer #1

we can equate the gravitational force from both planets

GMam / ra2 = GMbm / rb2

we get

Ma / ra2 = Mb / rb2

ra / rb = sqrt (ma / mb)

ra / rb = sqrt (0.6mb / mb)

ra / rb = sqrt (0.6)

ra / rb = 0.7745